OpenStudy (anonymous):

simplify completely 2x^2+20x+32 ________________ x^2-2x-80

OpenStudy (anonymous):

numerator: 2x^2+20x+32 =2(x^2 +10x +16) =2(x^2 + 8x + 2x +16) =2(x(x+8)+2(x+8)) 2(x+2)(x+8) denominator: x^2-2x-80 =x^2 -10x + 8x -80 =x(x-10)+ 8(x-10) =(x-10)(x+8) so, 2(x+2)(x+8) ------------ (x-10)(x+8) = 2(x+2) ------- (x-10)